A Complete Bibliography of Publications in the Aleph: Historical Studies in Science and

Judaism

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

29 February 2024 Version 1.00

Title word cross-reference

⊃ [Lan06]. Δ [Hal13].

/order [FF12b].

117b [BHKAM19]. 15th [Tau10]. 15th-Century [Tau10]. 1917-2000 [Lan01b]. 1920s [Pea15].

2003 [Rei03]. 247-Year [Vid17b, Vid17a].

535 [SI15].

īliyyāt [Lan06].

Abū [Sel22b]. Abraham [Ano17c, Arr11, BHM21, Fre16, Gar20, Gol11a, Gol20, Lév01, LB06, Rob21, SF06, Sel09, SS09, Sel10, Sel13, Sel17, Sel19, Sel20, Sel22a, Sel22b, Smi06, SL16, Vis22, ZF09]. Abravanel [Eli12]. Abstracts [Ano04a]. Access [Man12]. According [Ara06, Arr11]. Account [Gaz12, Rav08]. Accounts [KB10]. Adaptation [KS10]. Added [Eli18]. Ages [Men09]. al [Dav14, Vid21, Sel16, BK16, Fen16, Lan11, Lév02, Rob03, Rob09]. Al-Farghānī [Sel16]. Albert [Rei07]. Albertus [Lan05]. Alchemical [Lan05, Lan13a]. Alchemist [ZP20]. Alchemists [Men09]. Alexandria [LV18]. Alfonsine [Gol19a]. Alfonso [Gla13, Val10]. Algebra [Lév02]. alias [Mor11]. alongside [Sha18]. Alphabet [Lan03, Lan06]. Alter [Fre11a, FR11]. Alternative [Lév12]. American [Pea15]. Amos [Ano03a]. analysis [BHM21]. Analytics [Lan17b]. Ancient [Eli05, Fun03, Ste05, Lan06]. Andalusian [Del08]. Animals [Lan11, Gaz12]. Anniversary [Sel09, Sel13]. Annotated [RAK20]. Anonymous [Ara13, Gla09]. Answers [KS12]. Antecedents [BK16]. Appendix [Bor17, Fre17, Lan17a]. Application [KB16]. Approach [Xiu22]. Appropriations [Mes14]. April [Rei03]. Arabic [Asl02, BG07, BM15, BK16, BK22, But10, FV09, Lan14, Lan19, SL16]. Aramaic [Boh18]. Archimedean [Lév12]. Archimedes [Gla13]. Area [Lév12]. ’ares [Rub08]. Argument [Gla01]. Aristotelian [Eis21, KB16]. Aristotle [Eis21, Fon01, Hal13, KS10, KB10, KBS12, Lan16, Mey22, Rav08, Zon17]. Arithmetic [Ara13, LB06]. Asaf [YT18]. Ashkenazi [Lei13]. Astral [Sch03]. Astrolabe [RAK20]. Astrological [BS04, Gar20, KS18, Sel14, Sel19, Sel20, Smi06, SI15]. Astrological-Astronomical [Sel14]. Astrology [Ano17c, Arr11, Fre19, Lei13, Sel01, Sel04, Sel12b, Sel17]. Astronomer [Alt11, FR11]. astronomiae [Man17]. Astronomical [Bur17, Gar20, Gol05, Gol13a, GC17, Gol22, Mor11, Sel13, Sel14, SM07]. Astronomical-Astrological [Gar20]. Astronomy [FR11, Gol01, Gol10, Gol20, GH20, Man12, Nor02, Rob03, Sel01, Val10, MF05]. Attias [Bre07]. Attitudes [Las10]. Attributed [Lan05, Lan13a, LB06]. August [Rei03]. Augustinus [GC21a, GC21b]. Autograph [Lac03]. Avendauth [Fre16, Szi16]. Averroe [HH21, Har12a]. Averroes [Eis21, Fra08, Gla01, Hal13, HH21, Mey22, Zon17]. Avicennan [Lan12]. Avicennian [Fre08c]. Avigdor [BHM21]. Await [Lau02]. Azareus [Mes14].

B [Lan17a]. Back [Lau02]. Bar [Arr11, Sel13, ZF09]. Barnabas [Gla18]. Barren [Lan12]. Baruch [Kog09]. Bate [Sel20]. be [Ran09]. be-rabbi [Ran09]. Beginnings [Lan02b]. Ben [Man02, Fre19, Gar20, Gol02, Gol10, Gol11b, Gol12, GC17, Gol17, Gol20, GH20, GC21b, Gol22, Lan09, Lév12, MF05, Man12, Man17, RAK20, Vis14, Fre18b]. Ben-Shalom [Fre18b]. Berechiah [Vis14]. Bereshit [Ran09]. Between [MC02, Coh07, Kog09, LV18, Sel01]. Beyond [Har18]. Béziers [Bur17]. Bible [BS11]. Biblical [But10, KB10, KB16]. Bibliographia [Kel03]. Bibliography [Ano03a]. bin [Lan04b]. Bird [But10]. Bisection [Gla01]. Bitrūjī’s [Rob03]. Bocarro [MC02]. Bodleian [BHKAM19]. Bonfils [Gol17]. Book [FV09, FS17, Lan17b, Szi16, YT18, Gaz12]. Books [Ano01d, Ano02b, Ano03c, Ano04c, Ano05b, Ano07b, Ano09c, Ano10c, Ano11c, Ano13c, Ano15c, Ano18, BG16]. Borders [Sel01]. Botanical [BM15, BHKAM19]. Botany [Cha03]. Breggren [Gla18]. Bridge [Coh07]. Bright [Gol18b]. Budapest [Szi16]. Byzantine [Sch03].

C [Fre17]. Cairo [Boh18, BHM21]. Calculation [BS04]. Calendar [SM07, SM08, SI15, Vid17a, Vid17b, Vid22]. Calendars [Eli05, Ste05, Vid21]. Case [Eli18, GH20, Isr04, Kog09]. Categories [Rob09]. Center [Bur17]. Central [Men09]. Centuries [Rud02]. Century [BS11, Che03, Fon01, Fre16, Fre18a, Isr04, Sha18, Tau10, Bre07, Bur17, FM14, GC15, Hal13, Lan04b, LB06, Lév12, LV18, MC02, Sel01, SM07, Vis14, Zal05]. Chance [KBS12, KB16]. Chapter [Lan10]. Characters [BHKAM19]. Chelm [Rub08]. Chemical [Coh07]. China [Lan19]. Christian [Gla13, BS11, Lan04b]. Christians [Bur17]. Chronological [SF06]. Church [Fre18b]. Circa [ZF09]. Circle [Lév12, Sch01]. Citation [Lan04a]. Cities [Gol18a]. Class [Kra08]. Classical [Fun03]. Climates [AS05]. Clouds [Ran09]. Cod [BK22]. Collection [Szi16, Lan16]. Colors [Gol05]. Commentaries [Gla09, Lan16, Zon01]. Commentary [Hal13, Lan04a, Rob21, Sel04, Sel13, Zon12, Zon17, ZP20, Sha18]. Commentator [HH21]. Commenting [HH21]. Community [Gol01]. Comparison [Lan13b]. Compendium [Gla19, Lan07]. Composition [SL16]. Conatus [Har12b]. Concept [KBS12, Str01]. Conceptions [Sch03]. Concepts [Shi04]. Conclusions [Lau02]. Concordance [BS11]. Connection [Sel19, Shu02]. Constantinople [LV18]. Constellations [Sel16]. Construct [Eli18]. Construction [RAK20]. Contacts [MC02]. Context [BS11, Del08, Slá22]. Contexts [Koz11]. Contribution [Val10]. Contributions [Nor02]. Contributors [Ano01a, Ano02a, Ano03b, Ano04b, Ano05a, Ano06, Ano07a, Ano08, Ano09a, Ano09b, Ano10a, Ano10b, Ano11a, Ano11b, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, Ano16a, Ano16b, Ano17a, Ano17b]. Controversial [Bla08]. Coordinates [Gol18a]. Corporeal [Eli12]. Corpus [LV18]. Corrigenda [SM08]. Cosmogony [Xiu22]. Cosmogony-Prophetology [Xiu22]. Cosmology [Gol12, GH20]. Creation [Har18, Rav08, Vis21, Vis22]. Crescas [Ack17, Fra09, Lan12]. Critical [FS17, Har08, Sel04, Sha18, Zon03]. Criticism [Eis21, MF05]. Critique [Ste05]. Cross [Gol11b]. Cultural [Zal05]. Culture [Lei13]. Cycle [Vid17a, Vid17b].

Daily [GC15]. Dan [Sha18]. Daniel [BS04]. Darwinism [Lan13b, Pea15].

Data [Zon01]. Daud [Fre16, Lan16, Szi16, SL16]. David [Alt11, Fre11a, Lan16, Slá22, Szi16, Slá22]. Davke [Ber09]. Days [BS04]. Debate [Fel12]. Dedicatory [Man17]. Definitions [FS17]. Delmedigo [Lan13a]. Derrida [Har18]. Diaspora [Vid22]. Dictionary [Fen16]. Did [Asl02]. Differences [Pea15]. Diplomatic [RAK20]. Dis [FF12b]. Dis- [FF12b]. Dis-/order [FF12b]. Discovered [Gla01, Lév02, MC02, Sel16]. Discoveries [Smi06]. Disenchantment [Fun03]. Dodi [Vis14]. Dominicus [Fre16]. Donnolo [SM08, SM07]. Dossier [Fre11a]. Duhem [Fel12]. Duhem-Pines [Fel12]. Dynamics [Lan12].

Earliest [SI15]. Early [Dav14, Rub08, Rud02, But10, Eli21, Gol01, Lan04b]. East [Zal05]. Eastern [Las10, Tau10]. Eclipses [Gol05]. ed [Sha18]. Edition [Eli21, RAK20, Sel04, Sel13, Sel22b, Sha18, SL16, Zon03, BHM21]. Editions [FS17]. Editor [FF12a, FL22, LF19, Fre01, Fre05, Fre06, Fre08a, Fre09, Fre11b, Fre12]. Education [Rei07]. Eighteenth [Rud02, Bre07]. Eighteenth-Century [Bre07]. Eighth [GC21a]. Einstein [Rei07]. Eisagoge [Rob09]. Elder [Lan17b]. Elea [Gla01]. Eleazar [Ran09]. elements [BHM21, Eli18, Eli21]. Elior [Ste05]. Elusive [ZP20]. Embedded [Sel13]. Emergence [Fun03]. Empire [Sch03]. Encounter [LV18]. Encyclopedia [Sel14]. End [BS04]. English [Sha18, Sel22b]. Engraved [Mes14]. Epistemological [Fre19]. Epistemology [Har08]. Epitome [Eis21, Mey22]. Epitomes [Bos09]. Erratum [Sel12a, Vid17b]. Error [Gol10]. Essay [Sel13]. Eternal [Fre08c]. Euclid [Eli18, Eli21]. Europe [Men09]. European [Las10, Zal05, Rud02]. Evidence [Gla01]. Evolution [Pea15, Sch01]. Evolutionary [Che03]. ex [HH21, SI15]. ex-Sassoon [SI15]. Example [Hal13]. Exegesis [KB16]. Exegetical [Sel19]. Exhaustive [Smi06]. Existence [Ack17]. Experience [Lan14]. Explanations [Ara06]. Exploring [YT18]. Extensa [Fra09]. Ezra [Ano17c, Ara06, Arr11, Lév01, LB06, Rob21, SF06, Sel09, SS09, Sel10, Sel17, Sel19, Sel20, Sel22a, Sel22b, Smi06, Vis22].

Fārābī’s [Rob09]. Farghānī [Sel16]. Fevers [FV09, Lan17b]. Fifteenth [BS11, Sha18, GC15, LV18]. Fifteenth-Century [Sha18, GC15, LV18]. Finzi [Lac03]. First [BS11, Rob03]. fl [Gol20]. fol [BHKAM19]. Folkters [Gla18]. Foreword [FF12a, FL22, Fre01, Fre05, Fre06, Fre09, Fre11b, Fre12, LF19, Fre08a]. Forged [ZP20]. Form [Eli12]. Found [Fen16]. Foundations [Fre19]. Fourteenth [Fre18a, Hal13, Lév12]. Fourteenth-Century [Fre18a]. Fragment [Gol13b, Lan08a, Sel10, Szi16, SL16, BHM21]. Fragments [Lan16, SS09]. Francês [MC02]. France [Sel04]. Franco [Gol13a]. Franz [Rei03]. French [FM14, Sel22b]. Funkenstein [Ano03a]. Further [Was02]. Fuzzy [Sel01].

Galeano [Mor11, Lan07, Lan09]. Galenic [BG07, Bos09]. Galilei [MC02].

Galileo [MC02]. Ganāh [BK16, Fen16]. Gans [Alt11, Fre11a, Slá22]. Gaon [Shi04, Vid21, Vis21]. Geminos [Lan10]. General [Gar20]. Genesis [Lan04a]. Geniza [Lan08a]. Genizah [Lan16, Szi16, Boh18, BHM21]. Geographical [Gol18a, Rub08]. Geometrical [Gla19]. Geometry [LB06, Lév12]. Georg [Fre11a, FR11]. Gerard [BHM21]. Gerasa [ZF09]. German [Kog09]. Germany [Ano01c, Vol01]. Gershom [Fre19, MF05]. Gershom’s [Man17]. Gerson [Gol02, Gol10, Gol11b, Gol12, GH20, GC21b, Man02, Man12]. Gersonideana [Kel03]. Gersonides [Ano17c, Eis21, FF12b, Gaz12, Gla02, Har12b, KB10, KBS12, KB16, Sel17]. Gersonides. [Fre18a]. Glasner [Fre18a, Har12a]. Glossaries [BM15]. Glossary [BHKAM19]. Glosses [Zon17]. God [Fra09]. Greece [Fun03]. Greek [LV18]. Guide [Bla08, Dav08, Fre08b, Har08, Ivr08]. Gundissalinus [Fre16].

ha-’ares [Rub08]. ha-Middot [LB06]. ha-Millot [Rob09]. ha-Naqdan [Vis14]. ha-Se’elot [SS09]. ha-Tequfah [Sel09]. ha-Zarot [Rob09]. Hagin [Sel22b]. Hakam [Lan04b]. Halevi [AS05]. Hans [Lan13b]. Hasdai [Ack17, Fra09, Lan12]. Haskalah [Kog09]. Heavens [Fre08b, Rob21]. hebr [BK22]. Hebrew [Sha18, Ara13, BS11, BM05, BG07, BM15, BHKAM19, Del08, Eli18, Eli21, FV09, Fon01, FM14, FS17, Gla01, Gla09, Gla13, Gla19, Gol05, Gol13b, GC15, Gol18a, Gol18b, Gol19a, Gol19b, Gol20, Hal13, KS10, Koz11, Lan02b, Lan03, Lan04a, Lan05, Lan06, Lan10, Lan14, Lév01, Lév02, LV18, Mey22, Rob03, Sel04, SS09, Sel13, Sel16, Smi06, Vis14, YT18, Zon03, Zon17]. Henry [Sel20]. Hero [LV18]. Higher [Rei07]. Hippocrates [Zon03]. Historian [FR11]. Historical [Arr11, Zon03]. History [Cha03, Fre19, Lan02b]. Hitherto [Sel22b]. Hiyya [Arr11, Sel13, ZF09]. Hokmah [Lan09, Sel10, Lan07]. Hold [Lau02]. Holocaust [Lan13b]. Horologion [Boh18]. Horoscopy [Sel09, Sel13]. Hug [Ano17c, Rub08]. Hughes [Gla18]. Humankind [Fre19]. Hunayn [BG07].

Iberian [BM15, Nor02]. Ibn [Ano17c, Arr11, BG07, Fen16, Fra08, Fre08c, Fre16, KS12, Lév01, LB06, Rob09, Rob21, SF06, Sel09, SS09, Sel10, Sel17, Sel19, Sel20, Sel22a, Sel22b, Smi06, SL16, Vis22, BK16, Ara06, Dav14, Lan16, Szi16]. Ideal [Fun03]. Identification [BM05, Gla09]. Identified [Gol22, Sel13]. Identity [Szi16, ZP20]. Ignorant [Lau02]. II [Ano01c, Vol01]. Illustrated [BHKAM19]. Immanuel [Gol17, Gol22, Lév12, Rob21]. Imperial [Ano01c, Vol01]. Implicit [Zon12]. including [Lan11]. Indian [Shu02]. instrumenti [Man17]. Intellectuals [Sel01]. Interaction [Kog09]. Interest [Cha03]. Interesting [Lan17b]. Interpretation [Eis21, Vis21, Vis22]. Interpreters [Sch01]. Interpreting [Rav08]. Introduction [BS04, Gol17, KS18, Sha18, Zon03]. Investigative [KBS12]. Isaac [Eli12, FV09, FS17, GC17, Lan17b]. Ishāq [BG07]. Islam [Gol01]. Islamic [Lan06]. Isosceles [Eli18]. Isrā [Lan06]. Israel [Fre19, Fun03]. Israeli [Lan17b, FV09, FS17, GC17]. Istanbul [Gol20]. Italo [BHKAM19]. Italy [BM15, Isr04, SM07].

J [Gla18]. Jālīnūs [Mor11]. jâza [Ivr08]. Jacob [Gol17, Gol22, Lév12, MC02, Vid21]. Jakob [Cha03]. Jewish [Boh18, Alt11, BS11, BK16, Bre07, Cha03, Che03, Eli05, Fel12, Gol01, Isr04, Lan04b, Lan13b, Men09, Mes14, MC02, Nor02, Pea15, Rei07, RF15, Rud02, Sch03, Sel01, Ste05, SM07, SM08, Vid21, Vol01, Zal05, ZP20]. Jewry [Rud02]. Jews [Fre18b, Ano01c, Bur17, Hol02]. Jirí [FR11]. John [Gol13b, Vis21, Vis22]. Jonah [BK16, Fen16]. Jonas [Lan13b]. Joseph [Asl02, Bre07, GC17, Lan13a, RAK20]. Journal [Ber09]. Judah [Lan09, AS05, Zon01, Zon12]. Judaizers [Sha18]. Judeo [But10, Lan19, BK22, Lan14, SL16]. Judeo-Arabic [Lan19, Lan14]. Judgment [Lau02]. Juif [Sel22b]. Justification [Sch01].

Kafah [Lan01b]. Kaplan [Lan13b]. Karaite [Las10]. Kaspi [Asl02]. Katz [Gla18]. Kaufmann [Lan16, Szi16]. Key [Ber09]. Khwārizmī’s [Lév02]. Kitāb [Fen16, BK16]. Knaan [Tau10]. Knot [Ste08]. Know [Asl02]. Knowability [Fre08b]. Knowing [Lau02]. Knowledge [Fun03, Gla02, Vis14]. Komtino [LV18].

Languages [Gla02]. Languedoc [Fre18b]. Lapidary [FM14, Mes14]. Late [Rud02]. Later [Men09]. Latin [BM05, BG07, BHKAM19, FV09, KS10, Lév01, Smi06, Vis14]. Latitudes [Gol02]. Legacy [Vis22]. Lennart [Gla18]. Leon [Zon01, Zon12]. Letter [Man17, MC02]. Letters [Lan03, Bre07]. Levi [Fre19, Gar20, Gol02, Gol10, Gol11b, Gol12, GH20, GC21b, Man02, MF05, Man12, Man17]. Lexicography [Del08]. Library [Gol13b]. Light [Gaz12]. Limmudim [Kog09]. Limoges [Sel19]. Lindau [Kog09]. List [Man02, BK22]. Listing [SF06, Smi06]. Lists [Gol18a]. Literature [BM05, BK16, Coh07, Lan02b, Zal05]. Livorno [Bre07]. Livyat [Gar20]. Logic [Str14]. Logika [Sha18]. Long [Hal13, Lan04a, Lan11]. Lord [KBS12]. Lore [YT18]. Lost [Fen16].

Māshā’allāh [Sel12b]. Mūsā [Mor11]. Madda [RF15]. Magen [Slá22]. Magic [Lan09, Sch01, Sch03]. Magica [Lan11]. Magical [BS04, Sch01]. Magical-Talismanic [Sch01]. Magnus [Lan05]. Mahzor [SI15]. Maimonidean [Rav08]. Maimonides [Bos09, Dav14, Del08, Fra08, Fre08b, Har08, Sel04, Sel12b, Str01, Str14, Vis22, Xiu22]. Mandarins [Ano01c]. Manoel [MC02]. Manuscript [BHKAM19, GC15, Lac03, SI15]. Manuscripts [BM15]. Maqbilot [Lan02b]. Marbode [FM14]. Mardochaeus [ZP20]. Marginalia [Lan19]. Marwān [BK16]. Mas‘ūdī [Lan11]. Ma‘shar [Sel22b]. Masih [Lan04b]. Materia [Lan11, Lan19].

Mathematical [Lév01, Rob21]. Mathematicians [Lau02]. Mathematics [Fre16, LV18]. Matthias [Cha03]. Meaning [Ara06]. Measures [Bos09, BK22]. Mechanics [Lan09]. Mechanism [Gar20]. Medica [Lan11, Lan19]. Medical [BM05, Fen16, Koz11, Lan06]. Medicinal [Bos09, BK22]. Medicine [Lan09]. Medico [BM15, BHKAM19]. Medico-Botanical [BM15, BHKAM19]. Medieval [Fra08, GH20, Mes14, Ara13, Eli18, Fel12, FM14, Gla13, Gol05, Gol19b, KS10, Koz11, Lei13, Vid22]. Me’ir [BS11]. Memoriam [Rei03]. Mendelssohn [Lau02]. Menso [Gla18]. Messer [Zon01, Zon12]. Metaphysics [Hal13, Zon17]. Meteorological [Vis21, Vis22]. Meteorology [Fon01, KB10, Rav08]. Method [HH21]. Methodology [Cha03, GH20]. Metrological [LV18]. Microscopic [Cha03]. Mid [Bur17, Fre16, LB06]. Mid-Twelfth [Bur17, Fre16]. Mid-Twelfth-Century [LB06]. Middle [Men09, Zon17]. Middot [LB06]. Millot [Rob09]. Miracles [KB10]. Mistaken [ZP20]. Mivharim [SS09]. Modern [Eli21, Las10, RF15]. Modernization [Rei07, Rud02]. Modes [Rav08]. Moon [GC15, GC21b]. Moons [Gol13b]. Mordecai [Lan13b]. Mordekay [Lac03]. Mordekhai [LV18]. Mori [Lan01b]. Moses [KS12, Lan07, Lan09, Lau02, Mor11, Vis22]. Motion [GC21a]. Mover [Fre08b]. Ms [BHKAM19, BK22, Sel14, SI15]. Much [Asl02]. Multiple [Ack17]. Munich [BK22]. Musical [Shi04]. Musician [FR11]. My [Lan08b, Lan10, Lan14, Lan17b, Lan01a, Lan02a, Lan03, Lan04b, Lan05, Lan07, Lan08a, Lan09]. Mystery [Gol11a].

Nahmanides [Sch01]. Names [But10]. Naqdan [Vis14]. Narrative [Vis21, Vis22]. Nativ [BS11]. Natronai [Vis14]. Natural [Fre19]. Naturalistic [Gaz12]. Nature [Ber09]. Nazism [Rei07]. Nekdi [Vis14]. Nelle [ZP20]. Never [Ste08]. Newly [Gla01, Gol22, Lév02, MC02, Sel13, Sel16]. Newly-Discovered [Lév02]. Niccolò [Eli21]. Nicomachus [ZF09]. Nine [Zon17]. Nineteenth [Rud02, Zal05]. Nineteenth-Century [Zal05]. Ninth [Lan04b, Sel12b]. Nissayon [Lan14]. No [Lan12]. Non [Vid21]. Non-Rabbanite [Vid21]. Note [Ano07a, Ano08, Ano09a]. Notebooks [Lan01a, Lan02a, Lan03, Lan04b, Lan05, Lan07, Lan08a, Lan09, Lan10, Lan14, Lan17b]. Notes [Ano01a, Ano02a, Ano03b, Ano04b, Ano05a, Ano06, Ano09b, Ano10a, Ano10b, Ano11a, Ano11b, Ano12a, Ano12b, Ano13a, Ano13b, Ano14a, Ano14b, Ano15a, Ano15b, Ano16a, Ano16b, Ano17a, Ano17b, Gla09, Gol18a]. Notion [Fra09]. Number [Rob21].

Object [Gol18b]. Occitan [BHM21]. Old [BHM21, FM14]. Open [Fun03]. Opp [BHKAM19]. Optics [Lan08a]. Orb [Sel12b]. order [FF12b]. Origins [Ano01b, Vid17a, Vid17b, Was02]. Orthodox [Pea15]. Other [Lau02]. Overview [RF15]. Oxford [BHKAM19].

Page [BHKAM19]. Palestinian [Boh18]. Paracelsus [ZP20]. Parallels [Lan02b]. Paris [Sel14]. Part [Ano01c, Ano01b]. Partial [Lév02]. Parts [Vol01]. Passage [Dav08, Lan11]. Pedro [Val10]. Peninsula [BM15]. Perplexed [Dav08]. Perplexities [Lan08b]. Persian [YT18]. Perspective [Bla08, Rud02]. Perus [Rob09]. Peter [Sel19]. Pharmacognostic [BK16]. Philanthropismus [Kog09]. Philologist [Bla08]. Philoponus [Vis21, Vis22]. Philosopher [Fre18a]. Philosophers [Ack17]. Philosophical [Hal13]. Philosophy [Ber09, Fel12]. Physical [Fre19]. Physician [Lan04b]. Physics [Eis21, Gla09, KS12, Szi16, SL16, Zon01, Zon12, Lan16, Mey22, Har12a]. Pines [Fel12]. Piyyutim [Ran09]. Place [Fra09]. Planetary [Boh18, Gol02]. Plato [Har18]. Plural [Fel12]. Poetry [Koz11]. Polemics [BS11]. Portrait [Fre18a]. Positions [GC15]. Possible [Ack17]. Posterior [Lan17b]. posthumous [FS17]. Postscript [Ano17c]. Practica [BHM21]. Practical [Fre16]. Preeminent [Hol02]. Preliminary [Gol02, Gol13a, Gol19b]. Presented [Sha18]. Preserved [Gol13b]. Prevention [Rud02]. Principles [Gol12, Rob03]. Problematic [Dav08]. Problems [KS12]. Prognostication [KS18]. Prologue [Man17]. Proof [Lév12]. Prophecies [ZP20]. Prophetic [BS04]. Prophetology [Xiu22]. Provence [Fre18b]. Providence [FF12b, Fre19, Gar20, Gaz12]. Pseudo [BG07, Str01]. Pseudo-Galenic [BG07]. Pseudo-Science [Str01]. Ptolemaic [Sel16]. Ptolemy [MF05]. publication [FS17]. Published [Ano01d, Ano02b, Ano03a, Ano03c, Ano04c, Ano05b, Ano07b, Ano09c, Ano10c, Ano11c, Ano13c, Ano15c, Ano18, BG16]. Puzzle [Xiu22].

Qāfih [Lan01b]. Qaraite [Vid22]. Qiftī [Dav14]. Qillir [Ran09]. Qirqisānī [Vid21]. Qohelet [Ara06, Rob21]. Queries [KS12, Sel04]. Question [Isr04].

Rabbah [Ran09]. Rabbanite [Vid21]. Rabbi [BK16, Lan01b, Ran09, Rub08]. Rachel [Ste05]. Rain [Ran09]. Ram [Fre18b]. Rashi [Har18]. Rationalism [Sch03]. Ravings [Str01]. Reaction [Lan12]. Reagent [Lan12]. Recently [Ano01d, Ano02b, Ano03c, Ano04c, Ano05b, Ano07b, Ano09c, Ano10c, Ano11c, Ano13c, Ano15c, Ano18, BG16]. Recently-Published [Ano01d]. Reception [Fon01, Lei13]. Recess [Man12]. Reconsidered [Hol02]. recta [Lan05]. Redaction [Sel10]. Redactions [SS09]. Redemption [Fre19]. Rediscovered [Lac03]. References [Rob03]. Reflections [Har12a]. Refugees [Rei07]. regard [BHM21]. Regarding [Dav14]. Regimen [BG07]. Rejection [Eli12]. Religion [RF15]. Religious [Lan13b]. Remarks [Gol02, Gol13a, Gol19b, Lan17b]. Renaissance [Fre18b, Alt11, Lan07]. Rendering [But10]. Report [Gol18b]. Republic [Ano01c, Bre07]. Res [Fra09]. Re’sit [Sel10, Kog09]. Response [Eli05]. Responses [Che03, Pea15, Rud02]. review [Fre18a, Fre18b, Gla18, Sha18]. Revision [Hal13, Mey22]. Revisited [Gol11b]. Revisiting [Xiu22]. Ricius [GC21a, GC21b]. Rodez [KS18]. Roi [Gla18]. Romance [BM05, BM15, BHKAM19]. Rome [Rob21]. Rosales [MC02]. Rosenthal [Rei03]. Ruth [Fre18a, Har12a]. Ruthenian [Sha18]. Rylands [Gol13b].

Saadia [Shi04, BS04, Lan04a, Vid21, Vis21]. Sacrifice [Sch01]. samayim [Ano17c]. Samuel [Fra08, Fre08c, Rob09]. Sassoon [SI15]. Schieiden [Cha03]. Scholarly [SF06, Sel22a]. Scholarship [Bre07, Hol02]. Scholastic [Ack17]. Science [Ano01b, Bre07, Cha03, Dav14, Fra08, Hol02, Isr04, Lan07, Las10, RF15, Str01]. Scientific [Ano01c, Ara06, Fon01, GH20, Lan02b, Tau10, Vis14, Zal05]. Scientist [MC02, Fre18a]. Scientists [Vol01]. Sebastianist [MC02]. Secret [Rob21]. Section [Gar20, SI15]. Secular [Coh07]. Se’elot [SS09]. Sefer [LB06, Sel09, SS09, Shu02, Was02]. Semita [Lan05]. Sent [Sel04]. Sentence [Bla08]. Seventeenth [MC02]. Seventeenth-Century [MC02]. Shabbetai [SM08, SM07]. Shadow [Fre18b]. Shalom [Fre18b]. Shapira [Sha18]. Shows [Isr04]. Shtetl [Coh07]. Signature [Gol11a]. Signs [Lan10, Sel12b]. Simultaneous [Ack17]. Skandalon [Fra08]. Sky [Gol18b]. Slavic [Sha18]. Sleep [KS10]. Smallpox [Rud02]. Social [Ano01b]. Society [Zal05]. Solo [BHM21]. Solomon [RAK20, Rub08, Gol13a, Lan13a]. Solved [Gol11a]. Some [Ano01d, Ano02b, Lan17b, Men09, Nor02, Rud02, KB10]. Soul [FS17]. Source [Kog09]. Sources [Eli21, Gol10, KS12, Man12, Men09, Rub08, Sha18, Zon12]. Southern [Sel04]. Spain [Gla13, ZF09]. Spanish [FV09]. special [BHM21]. Speculation [BS04]. Sphere [GC21a]. Spinoza [Fra09, Har12b]. Spirit [FS17]. Staff [Gol11b]. Standardization [Hal13]. Star [Man02]. Statement [Dav14]. Stones [Mes14]. Studies [Lan10]. Study [BS11, Dav14, Kog09, RAK20]. Studying [Lan02b]. Sublunar [FF12b]. Success [Ano01b]. Supercommentaries [Rob21]. Supercommentary [Eis21]. superfoetatione [Zon03]. Synonyms [BM05]. System [Eis21].

Ta‘alumot [Lan09, Lan07]. Table [Gol13b, Gol22, SM07]. Tables [Gol05, Gol13a, GC15, GC17, Gol17, Gol19a, Gol19b, Sel13]. Taitasaq [RAK20]. Tajriba [Lan14]. Tajriba/Nissayon [Lan14]. Talhīs [Fen16, BK16]. Talismanic [Sch01]. Tarascon [Gol22, Lév12]. Tartaglia [Eli21]. Techel [Mes14]. Techel/Azareus [Mes14]. Tenth [SM07]. Tenth-Century [SM07]. Tequfah [Sel09]. Terminology [Arr11, Hal13]. Terms [BM05, BHKAM19]. Testimony [ZF09]. Text [KS18, Kra08, Lév01, LB06, Sel04, Sel16, Slá22, SL16]. Texts [Lan14, Sha18, Tau10]. Textual [FV09]. Their [Fre08b, BM05, Lau02, Sha18]. Theory [AS05, Che03, Fre08c, Gol02, KB16, Man12]. There [Kra08, Shu02]. Thesis [Fel12, Hol02]. Theurgy [Sch01]. Thirteenth [Fon01]. Thought [Lan13b, RF15, Sel01]. Thoughts [Was02]. Three [Che03, GC15, Lan19, Sel01]. Tibbon [Fra08, Fre08c, KS12, Rob09]. Todros [HH21]. Todrosi [HH21]. Toledo [Fre16, GC17]. Tool [KBS12]. Torah [RF15]. Tov [Gol20]. Tractatus [Man17]. Tradition [Eli18]. Traditionalist [Xiu22]. Traditions [FV09, Lan06]. Transcribed [Lan06]. Transfer [Vis14]. Transformation [Zal05]. Translating [Gla13]. Translation [Gla01, KS18, Lan04a, Mey22, RAK20, Sel04, Sel13, Sel16, Sel22b, Sha18, Zon03, BHM21]. Translations [BG07, But10, Gla13]. Translator [Sel20]. Transmission [Lan12, Mey22, Tau10]. Treatise [Ara13, BG07, KS10, Lan05, Lan13a, Mor11, Rub08, Sel09, Sel22b]. Treatises [Lan03]. Treatment [Ber09]. Triangle [Eli18]. Trigonometric [Gol19b]. Truest [Lan08b]. Tsarfat [FM14]. Turkey [Rei07]. Twelfth [LB06, Bur17, FM14, Fre16, Sel01, Vis14]. Twelfth-Century [FM14, Sel01]. Twentieth [Che03, Isr04]. Twentieth-Century [Che03]. Twenty [Zon17]. Twenty-Nine [Zon17]. Two [Gla09, Lan03, SS09, Zon12].

u [RF15]. u-Madda [RF15]. Uncovering [HH21]. Universes [Fel12]. Unknown [Ara13, KS10, Lan11, Lév01, Sel09, SS09, Sel10, Sel22b]. Update [Sel22a]. Upper [Ran09]. Use [KB10]. Uses [Koz11].

Valladolid [Gla13]. Variations [Eli18]. ve [Vis14]. Veblen [Hol02]. Version [BG07, Eli21, Lan05, Lév02]. Versions [FM14, FS17, Lév01]. Victor [Gla18]. Violas [KS18]. Vitry [SI15]. vs [BS04].

Wagner [Gla18]. Wakefulness [KS10]. Wars [KBS12]. Was [Ste08]. Waters [Ran09]. Weather [Lan10]. Weights [Bos09, BK22]. Weimar [Ano01c]. Within [Sch03, Fre19]. Work [Sel01]. Works [Shi04, Smi06, Vid21]. World [Coh07, Fra09, Fre08c, FF12b, Lan19]. Worlds [Ack17]. Writings [Ano03a, Fon01, SF06, Sel20, Sel22a].

Year [Vid17b, Vid17a]. Yerushalmi [Gol20]. yesirah [Shu02, Was02]. Yiddish [Ber09, Coh07]. Yom [Gol20]. Yosef [Lan01b]. Yusuf [Lan01b].

Zacut [Gol11a]. Zarot [Rob09]. Zeno [Gla01].

References

Ackerman:2017:UCS

[Ack17] Ari Ackerman. H .asdai Crescas and scholastic philosophers on the possible existence of multiple simultaneous worlds. Aleph: Historical Studies in Science and Judaism, 17(1):139–154, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

           URL https://muse.jhu.edu/pub/3/article/652313.

Alter:2011:DGR

[Alt11] George Alter. David Gans: a Renaissance Jewish astronomer. Aleph: Historical Studies in Science and Judaism, 11(1):61–112, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/423509#
sub01.

Anonymous:2001:NC

[Ano01a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 1(??):345–347, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/406604/pdf.

Anonymous:2001:PSO

[Ano01b] Anonymous. Part i: The social origins of success in science (1987). Aleph: Historical Studies in Science and Judaism, 1(??):216–249, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/406597#
sub01.

Anonymous:2001:PIJ

[Ano01c] Anonymous. Part II: Jews as scientific “Mandarins” in Imperial Germany and in the Weimar Republic (1997). Aleph: Historical Studies in Science and Judaism, 1(??):250–281, ???? 2001. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/406597#sub02.

Anonymous:2001:SRP

[Ano01d] Anonymous. Some recently-published books. Aleph: Historical Studies in Science and Judaism, 1(??):341–343, ???? 2001. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/406603/pdf.

Anonymous:2002:NC

[Ano02a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 2(??):303–305, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/407519/pdf.

Anonymous:2002:SRP

[Ano02b] Anonymous. Some recently — published books. Aleph: Historical Studies in Science and Judaism, 2(??):191–302, ???? 2002. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/407518/pdf.

Anonymous:2003:BPW

[Ano03a] Anonymous. A bibliography of the published writings of Amos Funkenstein (1937–1995). Aleph: Historical Studies in Science and Judaism, 3(??):83–95, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407522/pdf.

Anonymous:2003:NC

[Ano03b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 3(??):411–413, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/407533/pdf.

Anonymous:2003:RPB

[Ano03c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 3(??):375–410, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/407532/pdf.

Anonymous:2004:A

[Ano04a] Anonymous. Abstracts. Aleph: Historical Studies in Science and Judaism, 4(??):7–9, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/176039.

Anonymous:2004:NC

[Ano04b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 4(??):331–332, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/176038.

Anonymous:2004:RPB

[Ano04c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 4(??):299–330, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/176041.

Anonymous:2005:NC

[Ano05a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 5(??):363–365, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/187935.

Anonymous:2005:RPB

[Ano05b] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 5(??):303–362, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/187943.

Anonymous:2006:NC

[Ano06] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 6(??):399–400, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/203922.

Anonymous:2007:NC

[Ano07a] Anonymous. Note on contributors. Aleph: Historical Studies in Science and Judaism, 7(??):365–370, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/221070.

Anonymous:2007:RPB

[Ano07b] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 7(??):319–363, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/221072.

Anonymous:2008:NC

[Ano08] Anonymous. Note on contributors. Aleph: Historical Studies in Science and Judaism, 8(??):401–404, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247255.

Anonymous:2009:NCa

[Ano09a] Anonymous. Note on contributors. Aleph: Historical Studies in Science and Judaism, 9(1):183–184, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/260447.

Anonymous:2009:NCb

[Ano09b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 9(2):1–4, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/317030.

Anonymous:2009:RPB

[Ano09c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 9(1):141–180, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/260446.

Anonymous:2010:NCa

[Ano10a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 10(1):1–2, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/376444.

Anonymous:2010:NCb

[Ano10b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 10(2):1–2, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/394664.

Anonymous:2010:RPB

[Ano10c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 10(1):169–210, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/376445.

Anonymous:2011:NCa

[Ano11a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 11(1):205–206, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/423506.

Anonymous:2011:NCb

[Ano11b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 11(2):1–2, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/453374.

Anonymous:2011:RPB

[Ano11c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 11(1):179–204, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/423514.

Anonymous:2012:NCa

[Ano12a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 12(1):1–2, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/472832.

Anonymous:2012:NCb

[Ano12b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 12(2):1–3, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/487986.

Anonymous:2013:NCa

[Ano13a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 13(1):i, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/505934.

Anonymous:2013:NCb

[Ano13b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 13(2):1–3, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/523717.

Anonymous:2013:RPB

[Ano13c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 13(1):139–174, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/505933.

Anonymous:2014:NCa

[Ano14a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 14(1):1–2, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/543892.

Anonymous:2014:NCb

[Ano14b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 14(2):177–178, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/554173.

Anonymous:2015:NCa

[Ano15a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 15(1):200, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/576355.

Anonymous:2015:NCb

[Ano15b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 15(2):1–2, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/590709.

Anonymous:2015:RPB

[Ano15c] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 15(1):177–198, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/576354.

Anonymous:2016:NCa

[Ano16a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 16(1):239–240, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/612505.

Anonymous:2016:NCb

[Ano16b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 16(2):437, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/627993.

Anonymous:2017:NCa

[Ano17a] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 17(1):177–178, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/652316.

Anonymous:2017:NCb

[Ano17b] Anonymous. Notes on contributors. Aleph: Historical Studies in Science and Judaism, 17(2):373–374, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/669427.

Anonymous:2017:PGU

[Ano17c] Anonymous. Postscript: Gersonides’ h.ug šamayim, astrology, and Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 17(2):324–333, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/669423#sub01.

Anonymous:2018:RPB

[Ano18] Anonymous. Recently published books. Aleph: Historical Studies in Science and Judaism, 18(2):311–317, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/706092.

Aranda:2006:MQA

[Ara06] Mariano Gomez Aranda. The meaning of qohelet according to Ibn Ezra’s scientific explanations. Aleph: Historical Studies in Science and Judaism, 6(??):339–370, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/203923.

Aradi:2013:UMH

[Ara13] Naomi Aradi. An unknown Medieval Hebrew anonymous treatise on arithmetic. Aleph: Historical Studies in Science and Judaism, 13(2):235–309, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/523715.

Arribas:2011:THA

[Arr11] Josefina Rodríguez Arribas. The terminology of historical astrology according to Abraham Bar H .iyya and Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 11(1):11–54, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/423508.

Altmann:2005:JHT

[AS05] Alexander Altmann and Lenn J. Schramm. Judah Halevi’s theory of climates. Aleph: Historical Studies in Science and Judaism, 5(??):215–246, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/187936.

Aslanov:2002:HMA

[Asl02] Cyril Aslanov. How much Arabic did Joseph kaspi know? Aleph: Historical Studies in Science and Judaism, 2(??):259–270, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407515/
pdf.

Berger:2009:PYK

[Ber09] Shlomo Berger. Philosophy in a Yiddish key: The journal Davke’s treatment of nature. Aleph: Historical Studies in Science and Judaism, 9(1):113–140, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/260445.

Bos:2007:PGT

[BG07] Gerrit Bos and Ivan Garofalo. A pseudo-Galenic treatise on regimen: The Hebrew and Latin translations from H .unayn Ibn Ish.āq’s Arabic version. Aleph: Historical Studies in Science and Judaism, 7(??):43–95, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/221067.

Bos:2016:RPB

[BG16] Gerrit Bos and Ivan Garofalo. Recently published books. Aleph: Historical Studies in Science and Judaism, 16(2):407–435, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/627992.

Bos:2019:GLI

[BHKAM19] Gerrit Bos, Sandra Hajek, Katrin Kogman-Appel, and Guido Mensching. A glossary of Latin and Italo–Romance medicobotanical terms in Hebrew characters on an illustrated manuscript page (ms. Oxford, Bodleian opp. 688, fol. 117b). Aleph: Historical Studies in Science and Judaism, 19(2):169–199, ???? 2019. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/747571.

Bos:2021:FAA

[BHM21] Gerrit Bos, Sandra Hajek, and Guido Mensching. A fragment of Abraham Avigdor’s translation of Gerard de Solo’s practica from the Cairo Genizah: edition and analysis with special regard to the Old Occitan elements. Aleph: Historical Studies in Science and Judaism, 21(2):309–357, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/846000.

Bos:2016:APL

[BK16] Gerrit Bos and Fabian Käs. Arabic pharmacognostic literature and its Jewish antecedents: Marwān ibn Ǧanā\begin{matrix}h \\ . \end{matrix} (Rabbi Jonah), Kitāb al–Talh̆īs.. Aleph: Historical Studies in Science and Ju-

daism, 16(1):144–229, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/612504.

Bos:2022:JAL

[BK22] Gerrit Bos and Fabian Käs. The Judeo–Arabic list of medicinal measures and weights of MS Munich, cod. hebr. 275. Aleph: Historical Studies in Science and Judaism, 22(1–2):157–182, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/874964.

Blau:2008:CSG

[Bla08] Joshua Blau. The controversial sentence of guide 2.24: a Philologist’s perspective. Aleph: Historical Studies in Science and Judaism, 8(??):159–161, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/247251.

Bos:2005:LHM

[BM05] Gerrit Bos and Guido Mensching. The literature of Hebrew medical synonyms: Romance and Latin terms and their identification. Aleph: Historical Studies in Science and Judaism, 5(??):169–211, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/187937.

Bos:2015:ARM

[BM15] Gerrit Bos and Guido Mensching. Arabic–Romance medicobotanical glossaries in Hebrew manuscripts from the Iberian Peninsula and Italy. Aleph: Historical Studies in Science and Judaism, 15(1):9–61, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/576357.

Bohak:2018:PJA

[Boh18] Gideon Bohak. A Palestinian–Jewish Aramaic planetary horologion from the Cairo Genizah. Aleph: Historical Studies in Science and Judaism, 18(1):7–26, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/693213.

Borisov:2017:A

[Bor17] Andrei Iakovlevič Borisov. Appendix a. Aleph: Historical Studies in Science and Judaism, 17(1):76–87, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/652311#sub01.

Bos:2009:MMM

[Bos09] Gerrit Bos. Maimonides on medicinal measures and weights, from his Galenic epitomes. Aleph: Historical Studies in Science and Judaism, 9(2):255–276, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/317031.

Bregoli:2007:JSS

[Bre07] Francesca Bregoli. Jewish scholarship, science, and the Republic of Letters: Joseph Attias in Eighteenth-Century Livorno. Aleph: Historical Studies in Science and Judaism, 7(??):97–181, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/221068.

Ben-Shammai:2004:SID

[BS04] Haggai Ben-Shammai. Saadia’s introduction to Daniel: Prophetic calculation of the end of days vs. astrological and magical speculation. Aleph: Historical Studies in Science and Judaism, 4(??): 11–86, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
176040.

Ben-Shalom:2011:MNF

[BS11] Ram Ben-Shalom. Me’ir nativ: The first Hebrew concordance of the Bible and Jewish Bible study in the fifteenth century, in the context of Jewish–Christian polemics. Aleph: Historical Studies in Science and Judaism, 11(2):289–364, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/453376.

Burnett:2017:BAC

[Bur17] Charles Burnett. Béziers as an astronomical center for Jews and Christians in the Mid-Twelfth Century. Aleph: Historical Studies in Science and Judaism, 17(2):197–219, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/669421.

Butbul:2010:RBN

[But10] Sagit Butbul. The rendering of bird names in Early Judeo–Arabic biblical translations. Aleph: Historical Studies in Science and Judaism, 10(1):11–42, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/376439.

Charpa:2003:MJS

[Cha03] Ulrich Charpa. Matthias Jakob Schieiden (1804–1881): The history of Jewish interest in science and the methodology of microscopic botany. Aleph: Historical Studies in Science and Judaism, 3(??):213–245, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407526/pdf.

Cherry:2003:TTC

[Che03] Shai Cherry. Three Twentieth-century Jewish responses to evolutionary theory. Aleph: Historical Studies in Science and Judaism, 3(??):247–290, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407527/pdf.

Cohen:2007:CLY

[Coh07] Stephen M. Cohen. Chemical literature in Yiddish: a bridge between the shtetl and the secular world. Aleph: Historical Studies in Science and Judaism, 7(??):183–251, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/221069.

Davidson:2008:PPG

[Dav08] Herbert A. Davidson. The problematic passage in guide for the perplexed 2:24. Aleph: Historical Studies in Science and Judaism, 8(??):163–193, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/247263.

Davidson:2014:IAQ

[Dav14] Herbert A. Davidson. Ibn al–Qift.ī’s statement regarding Maimonides’ early study of science. Aleph: Historical Studies in Science and Judaism, 14(1):245–258, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/543890.

Delgado:2008:MCA

[Del08] José Martínez Delgado. Maimonides in the context of Andalusian Hebrew lexicography. Aleph: Historical Studies in Science and Judaism, 8(??):15–40, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/247262.

Eisenmann:2021:GCA

[Eis21] Esti Eisenmann. Gersonides’ criticism of the Aristotelian system and of Averroes’s interpretation of Aristotle in his supercommentary on the epitome of the physics. Aleph: Historical Studies in Science and Judaism, 21(1):79–121, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/799684.

Elior:2005:AJC

[Eli05] Rachel Elior. Ancient Jewish calendars: a response. Aleph: Historical Studies in Science and Judaism, 5(??):293–302, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/187938.

Elior:2012:IAR

[Eli12] Ofer Elior. Isaac Abravanel’s rejection of corporeal form. Aleph: Historical Studies in Science and Judaism, 12(2):367–402, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/487983.

Elior:2018:HCI

[Eli18] Ofer Elior. How to construct an isosceles triangle: Variations of a case added to Euclid’s Elements, in the Medieval Hebrew tradition. Aleph: Historical Studies in Science and Judaism, 18(2): 159–180, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
706084.

Elior:2021:NTE

[Eli21] Ofer Elior. Niccolò Tartaglia’s 1543 edition of Euclid’s Elements and the sources of an Early Modern Hebrew version of the Elements. Aleph: Historical Studies in Science and Judaism, 21(1): 123–148, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
799685.

Feldman:2012:PUD

[Fel12] Seymour Feldman. On plural universes: a debate in Medieval Jewish philosophy and the Duhem-pines thesis. Aleph: Historical Studies in Science and Judaism, 12(2):329–366, ???? 2012. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/487982.

Fenton:2016:JIU

[Fen16] Paul B. Fenton. Jonah Ibn Ǧanāh.’s medical dictionary, the Kitāb al–Talh̆īs.: lost and found. Aleph: Historical Studies in Science

and Judaism, 16(1):106–144, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/612503.

Fontaine:2012:EF

[FF12a] Resianne Fontaine and Gad Freudenthal. Editors’ foreword. Aleph: Historical Studies in Science and Judaism, 12(1):1, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/472824.

Freudenthal:2012:GDO

[FF12b] Gad Freudenthal and Resianne Fontaine. Gersonides on the dis- /order of the sublunar world and on Providence. Aleph: Historical Studies in Science and Judaism, 12(2):299–328, ???? 2012. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/487981.

Fontaine:2022:EF

[FL22] Resianne Fontaine and Reimund Leicht. Editors’ foreword. Aleph: Historical Studies in Science and Judaism, 22(1–2):1, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/874959.

Freudenthal:2014:OFH

[FM14] Gad Freudenthal and Jean-Marc Mandosio. Old French into Hebrew in Twelfth-Century Tsarfat: Medieval Hebrew versions of Marbode’s lapidary. Aleph: Historical Studies in Science and Judaism, 14(1):11–187, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/543888.

Fontaine:2001:RAM

[Fon01] Resianne Fontaine. The reception of Aristotle’s meteorology in Hebrew scientific writings of the thirteenth century. Aleph: Historical Studies in Science and Judaism, 1(??):101–139, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/406594/pdf.

Freudenthal:2011:GJA

[FR11] Gad Freudenthal and Jan Roubinek. Georg (Jiří) Alter (1891– 1972): Astronomer, historian of astronomy, and musician. Aleph: Historical Studies in Science and Judaism, 11(1):115–156, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/423509#
sub02.

Fraenkel:2008:MAS

[Fra08] Carlos Fraenkel. Maimonides, Averroes, and Samuel Ibn Tibbon on a skandalon of medieval science. Aleph: Historical Studies in Science and Judaism, 8(??):195–212, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247259.

Fraenkel:2009:UCG

[Fra09] Carlos Fraenkel. H .asdai Crescas on God as the place of the world and Spinoza’s notion of God as Res Extensa. Aleph: Historical Studies in Science and Judaism, 9(1):77–111, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/260444.

Freudenthal:2001:EF

[Fre01] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 1(??):9–10, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/406591/pdf.

Freudenthal:2005:EF

[Fre05] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 5(??):7, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/187939.

Freudenthal:2006:EF

[Fre06] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 6(??):7, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/203924.

Freudenthal:2008:EF

[Fre08a] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 8(??):13, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/247260.

Freudenthal:2008:MKH

[Fre08b] Gad Freudenthal. Maimonides on the knowability of the heavens and of their mover ( guide 2:24). Aleph: Historical Studies in Science and Judaism, 8(??):151–157, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247258.

Freudenthal:2008:SIT

[Fre08c] Gad Freudenthal. Samuel Ibn Tibbon’s Avicennian theory of an eternal world. Aleph: Historical Studies in Science and Judaism, 8(??):41–129, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/247261.

Freudenthal:2009:EF

[Fre09] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 9(1):11, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/260441.

Freudenthal:2011:DGA

[Fre11a] Gad Freudenthal. Dossier: Georg Alter (1891–1972) on David Gans (1541–1613). Aleph: Historical Studies in Science and Judaism, 11(1):56–59, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/423509.

Freudenthal:2011:EF

[Fre11b] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 11(1):8–10, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/423507.

Freudenthal:2012:EF

[Fre12] Gad Freudenthal. Editor’s foreword. Aleph: Historical Studies in Science and Judaism, 12(1):1, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/472823.

Freudenthal:2016:AID

[Fre16] Gad Freudenthal. Abraham Ibn Daud, Avendauth, Dominicus Gundissalinus and practical mathematics in Mid-Twelfth Century Toledo. Aleph: Historical Studies in Science and Judaism, 16(1): 60–106, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
612502.

Freudenthal:2017:AC

[Fre17] Gad Freudenthal. Appendix C. Aleph: Historical Studies in Science and Judaism, 17(1):90–93, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/652311#sub03.

Freudenthal:2018:GPF

[Fre18a] Gad Freudenthal. Gersonides. A Portrait of a Fourteenth-Century Philosopher–Scientist by Ruth Glasner (review). Aleph: Historical Studies in Science and Judaism, 18(2):283–287, ???? 2018. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/706089.

Freudenthal:2018:JPL

[Fre18b] Gad Freudenthal. (The Jews of Provence and Languedoc. Renaissance in the Shadow of the Church) by Ram Ben-Shalom (review). Aleph: Historical Studies in Science and Judaism, 18(2):288–294, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/706090.

Freudenthal:2019:PEF

[Fre19] Gad Freudenthal. The physical and epistemological foundations of Levi ben Gershom’s astrology: Providence and Israel’s redemption within the natural history of humankind. Aleph: Historical Studies in Science and Judaism, 19(1):59–130, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/724812.

Freudenthal:2017:HVI

[FS17] Gad Freudenthal and Samuel M. Stern. The Hebrew versions of Isaac Israeli’s book of definitions and book on spirit and soul: Critical editions (posthumous publication). Aleph: Historical Studies in Science and Judaism, 17(1):11–75, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/652311.

Funkenstein:2003:DKE

[Fun03] Amos Funkenstein. The disenchantment of knowledge: The emergence of the ideal of open knowledge in ancient Israel and in classical Greece. Aleph: Historical Studies in Science and Judaism, 3(??):15–81, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/407521/pdf.

Ferre:2009:TTI

[FV09] Lola Ferre and Raphaela Veit. The textual traditions of Isaac Israeli’s book on fevers in Arabic, Latin, Hebrew, and Spanish. Aleph: Historical Studies in Science and Judaism, 9(2):309–334, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/317029.

Garshtein:2020:MGP

[Gar20] Niran Garshtein. The mechanism of general Providence in the astronomical-astrological section of Levi ben Abraham’s livyat h.en. Aleph: Historical Studies in Science and Judaism, 20(1–2): 203–228, ???? 2020. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
776500.

Gaziel:2012:GNA

[Gaz12] Ahuva Gaziel. Gersonides’ naturalistic account of Providence in light of the Book of Animals. Aleph: Historical Studies in Science and Judaism, 12(2):243–271, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/487979.

Goldstein:2015:TTD

[GC15] Bernard R. Goldstein and José Chabás. Three tables for the daily positions of the Moon in a Fifteenth-Century Hebrew manuscript. Aleph: Historical Studies in Science and Judaism, 15(2):319–341, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/590708.

Goldstein:2017:ATI

[GC17] Bernard R. Goldstein and José Chabás. The astronomical tables of Isaac ben Joseph Israeli of Toledo. Aleph: Historical Studies in Science and Judaism, 17(2):357–370, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/669425.

Goldstein:2021:ARM

[GC21a] Bernard R. Goldstein and José Chabás. Augustinus ricius, on the motion of the eighth sphere. Aleph: Historical Studies in Science and Judaism, 21(2):359–377, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/846001.

Goldstein:2021:LBG

[GC21b] Bernard R. Goldstein and José Chabás. Levi ben Gerson and Augustinus Ricius on the Moon. Aleph: Historical Studies in Science and Judaism, 21(1):191–198, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/799682.

Goldstein:2020:SMM

[GH20] Bernard R. Goldstein and Giora Hon. Scientific methodology in medieval astronomy and cosmology: The case of Levi ben Gerson (1288–1344). Aleph: Historical Studies in Science and Judaism, 20(1–2):229–261, ???? 2020. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/776501.

Glasner:2001:ZEA

[Gla01] Ruth Glasner. Zeno of Elea’s argument from bisection: Newly discovered evidence in a Hebrew translation of Averroes. Aleph: Historical Studies in Science and Judaism, 1(??):283–293, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/406598/pdf.

Glasner:2002:GKL

[Gla02] Ruth Glasner. On Gersonides’ knowledge of languages. Aleph: Historical Studies in Science and Judaism, 2(??):235–257, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407514/
pdf.

Glasner:2009:TNI

[Gla09] Ruth Glasner. Two notes on the identification of two anonymous Hebrew commentaries on the physics. Aleph: Historical Studies in Science and Judaism, 9(2):335–344, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/317028.

Glasner:2013:HTM

[Gla13] Ruth Glasner. Hebrew translations in Medieval Christian Spain: Alfonso of Valladolid translating Archimedes? Aleph: Historical Studies in Science and Judaism, 13(2):185–199, ???? 2013. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/523716.

Glasner:2018:MFB

[Gla18] Ruth Glasner. Menso Folkters, Barnabas Hughes, Roi Wagner, and J. Lennart Breggren by Victor J. Katz (review). Aleph: Historical Studies in Science and Judaism, 18(2):279–283, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/706088.

Glasner:2019:HGC

[Gla19] Ruth Glasner. The Hebrew geometrical compendium. Aleph: Historical Studies in Science and Judaism, 19(2):201–268, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/747572.

Goldstein:2001:AJC

[Gol01] Bernard R. Goldstein. Astronomy and the Jewish community in Early Islam. Aleph: Historical Studies in Science and Judaism, 1(??):17–57, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/406592/pdf.

Goldstein:2002:LBG

[Gol02] Bernard R. Goldstein. Levi ben Gerson’s preliminary remarks for a theory of planetary latitudes. Aleph: Historical Studies in Science and Judaism, 2(??):15–27, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/407504/pdf.

Goldstein:2005:CEM

[Gol05] Bernard R. Goldstein. Colors of eclipses in Medieval Hebrew astronomical tables. Aleph: Historical Studies in Science and Judaism, 5(??):11–34, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/187940.

Goldstein:2010:LBG

[Gol10] Bernard R. Goldstein. Levi ben Gerson on the sources of error in astronomy. Aleph: Historical Studies in Science and Judaism, 10(2):211–240, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/394665.

Goldstein:2011:AZS

[Gol11a] Bernard R. Goldstein. Abraham Zacut’s signature: a mystery solved. Aleph: Historical Studies in Science and Judaism, 11(1): 159–167, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
423512.

Goldstein:2011:LBG

[Gol11b] Bernard R. Goldstein. Levi ben Gerson and the cross staff revisited. Aleph: Historical Studies in Science and Judaism, 11(2): 365–383, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
453377.

Goldstein:2012:LBG

[Gol12] Bernard R. Goldstein. Levi ben Gerson on the principles of cosmology. Aleph: Historical Studies in Science and Judaism, 12(1): 17–35, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
472825.

Goldstein:2013:PRA

[Gol13a] Bernard R. Goldstein. Preliminary remarks on the astronomical tables of Solomon Franco. Aleph: Historical Studies in Science and Judaism, 13(2):175–184, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/523714.

Goldstein:2013:TNM

[Gol13b] Bernard R. Goldstein. A table of new moons from 1501 to 1577 in a Hebrew fragment preserved in the John rylands library. Aleph: Historical Studies in Science and Judaism, 13(1):11–26, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/505929.

Goldstein:2017:IIB

[Gol17] Bernard R. Goldstein. The introduction to Immanuel ben Jacob Bonfils’ tables for 1340. Aleph: Historical Studies in Science and Judaism, 17(1):167–176, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/652315.

Goldstein:2018:NHL

[Gol18a] Bernard R. Goldstein. Notes on Hebrew lists of geographical coordinates of cities. Aleph: Historical Studies in Science and Judaism, 18(2):265–271, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/706086.

Goldstein:2018:RHB

[Gol18b] Bernard R. Goldstein. A report in Hebrew of a bright object in the sky in 1473. Aleph: Historical Studies in Science and Judaism, 18(2):273–278, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/706087.

Goldstein:2019:ATH

[Gol19a] Bernard R. Goldstein. The Alfonsine tables in Hebrew. Aleph: Historical Studies in Science and Judaism, 19(2):269–280, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/747573.

Goldstein:2019:PRM

[Gol19b] Bernard R. Goldstein. Preliminary remarks on Medieval Hebrew trigonometric tables. Aleph: Historical Studies in Science and Judaism, 19(1):131–136, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/724813.

Goldstein:2020:AHI

[Gol20] Bernard R. Goldstein. Astronomy in Hebrew in Istanbul: Abraham ben Yom T .ov Yerushalmi (fl. 1510). Aleph: Historical Studies in Science and Judaism, 20(1–2):303–307, ???? 2020. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/776496.

Goldstein:2022:NIA

[Gol22] Bernard R. Goldstein. A newly identified astronomical table by Immanuel ben Jacob of Tarascon. Aleph: Historical Studies in Science and Judaism, 22(1–2):279–285, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/874967.

Halper:2013:RSH

[Hal13] Yehuda Halper. Revision and standardization of Hebrew philosophical terminology in the Fourteenth Century: The example of Averroes’s long commentary on Aristotle’s Metaphysics Δ. Aleph: Historical Studies in Science and Judaism, 13(1):95–137, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/505932.

Harvey:2008:MCE

[Har08] Warren Zev Harvey. Maimonides’ critical epistemology and guide 2:24. Aleph: Historical Studies in Science and Judaism, 8(??): 213–235, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
247254.

Harvey:2012:RRG

[Har12a] Steven Harvey. Reflections on Ruth Glasner’s Averroes’ Physics. Aleph: Historical Studies in Science and Judaism, 12(2):403–412, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/487984.

Harvey:2012:GSC

[Har12b] Warren Zev Harvey. Gersonides and Spinoza on Conatus. Aleph: Historical Studies in Science and Judaism, 12(2):273–297, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/487980.

Harvey:2018:RCB

[Har18] Warren Zev Harvey. Rashi on creation: Beyond Plato and Derrida. Aleph: Historical Studies in Science and Judaism, 18(1): 27–49, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
693214.

Harvey:2021:AEA

[HH21] Steven Harvey and Oded Horezky. Averroes ex Averroe: Uncovering T .odros T .odrosi’s method of commenting on the commentator. Aleph: Historical Studies in Science and Judaism, 21(1):7–78, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/799683.

Hollinger:2002:WJP

[Hol02] David A. Hollinger. Why are Jews preeminent in science and scholarship?: The Veblen thesis reconsidered. Aleph: Historical Studies in Science and Judaism, 2(??):145–163, ???? 2002. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/407509/pdf.

Israel:2004:SJQ

[Isr04] Giorgio Israel. Science and the Jewish question in the Twentieth century: The case of Italy and what it shows. Aleph: Historical Studies in Science and Judaism, 4(??):191–261, ???? 2004. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/176043.

Ivry:2008:GAJ

[Ivr08] Alfred Ivry. Guide 2:24 and all that (i)jâza. Aleph: Historical Studies in Science and Judaism, 8(??):237–246, ???? 2008. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/247248.

Klein-Braslavy:2010:GUA

[KB10] Sara Klein-Braslavy. Gersonides’ use of Aristotle’s meteorology in his accounts of some biblical miracles. Aleph: Historical Studies in Science and Judaism, 10(2):241–313, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/394666.

Klein-Braslavy:2016:ATC

[KB16] Sara Klein-Braslavy. The Aristotelian theory of chance and its application in Gersonides’ biblical Exegesis. Aleph: Historical Studies in Science and Judaism, 16(2):367–405, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/627991.

Klein-Braslavy:2012:ACC

[KBS12] Sara Klein-Braslavy and Lenn Schramm. Aristotle’s concept of chance as an investigative tool in Gersonides’ Wars of the Lord. Aleph: Historical Studies in Science and Judaism, 12(1):65–100, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/472827.

Kellner:2003:BG

[Kel03] Menachem Kellner. Bibliographia Gersonideana, 1992–2002. Aleph: Historical Studies in Science and Judaism, 3(??):345–374, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407531/
pdf.

Kogman:2009:BLR

[Kog09] Tal Kogman. Baruch Lindau’s Rešit Limmudim (1788) and its German source: a case study of the interaction between the Haskalah and German philanthropismus. Aleph: Historical Studies in Science and Judaism, 9(2):277–304, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/317032.

Kozodoy:2011:MHM

[Koz11] Maud Kozodoy. Medieval Hebrew medical poetry: Uses and contexts. Aleph: Historical Studies in Science and Judaism, 11(2): 213–288, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
453375.

Kraemer:2008:TTC

[Kra08] Joel L. Kraemer. Is there a text in this class? Aleph: Historical Studies in Science and Judaism, 8(??):247–299, ???? 2008. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/247247.

Kahana-Smilansky:2010:ASW

[KS10] Hagar Kahana-Smilansky. Aristotle on sleep and wakefulness: a Medieval Hebrew adaptation of an unknown Latin treatise. Aleph: Historical Studies in Science and Judaism, 10(1):67–118, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/376441.

Kahana-Smilansky:2012:MIT

[KS12] Hagar Kahana-Smilansky. Moses Ibn Tibbon’s answers to queries on physics: Sources and problems. Aleph: Historical Studies in Science and Judaism, 12(2):209–241, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/487978.

Kahana-Smilansky:2018:VRA

[KS18] Hagar Kahana-Smilansky. Violas de Rodez’s astrological prognostication for 1355: Introduction, text and translation. Aleph: Historical Studies in Science and Judaism, 18(1):51–121, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/693215.

Lacerenza:2003:RAM

[Lac03] Giancarlo Lacerenza. A rediscovered autograph manuscript by mordekay finzi. Aleph: Historical Studies in Science and Judaism, 3(??):301–325, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407529/pdf.

Langermann:2001:MN

[Lan01a] Y. Tzvi Langermann. From my notebooks. Aleph: Historical Studies in Science and Judaism, 1(??):325–332, ???? 2001. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/406601/pdf.

Langermann:2001:MYR

[Lan01b] Y. Tzvi Langermann. “Mori yusuf”: Rabbi yosef Kafa\begin{matrix}h \\ . \end{matrix} (Qāfi\begin{matrix}h \\ . \end{matrix}) (1917-2000). Aleph: Historical Studies in Science and Judaism, 1(??):333–340, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/406602/pdf.

Langermann:2002:MN

[Lan02a] Y. Tzvi Langermann. From my notebooks. Aleph: Historical Studies in Science and Judaism, 2(??):279–290, ???? 2002. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/407517/pdf.

Langermann:2002:BHS

[Lan02b] Y. Tzvi Langermann. On the beginnings of Hebrew scientific literature and on studying history through “Maqbilot̄” (parallels). Aleph: Historical Studies in Science and Judaism, 2(??):165–189, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407510/
pdf.

Langermann:2003:MNT

[Lan03] Y. Tzvi Langermann. From my notebooks: Two treatises on the letters of the Hebrew alphabet. Aleph: Historical Studies in Science and Judaism, 3(??):293–299, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/407528/pdf.

Langermann:2004:CSL

[Lan04a] Y. Tzvi Langermann. A citation from Saadia’s long commentary to Genesis, in Hebrew translation. Aleph: Historical Studies in Science and Judaism, 4(??):293–297, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/176045.

Langermann:2004:MNM

[Lan04b] Y. Tzvi Langermann. From my notebooks: Masih bin Hakam, a Jewish–Christian (?) physician of the Early Ninth Century. Aleph: Historical Studies in Science and Judaism, 4(??):283–292, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/176044.

Langermann:2005:MNH

[Lan05] Y. Tzvi Langermann. From my notebooks: The Hebrew version of Semita recta, an alchemical treatise attributed to Albertus Magnus. Aleph: Historical Studies in Science and Judaism, 5(??): 273–284, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
187941.

Langermann:2006:MIA

[Lan06] Y. Tzvi Langermann. Medical isrā^{⊃}īliyyāt? Ancient Islamic medical traditions transcribed into the Hebrew alphabet. Aleph: Historical Studies in Science and Judaism, 6(??):373–398, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/203925.

Langermann:2007:MNC

[Lan07] Y. Tzvi Langermann. From my notebooks: a compendium of Renaissance science: Ta‘alumot H .ok̄mah by Moses Galeano. Aleph: Historical Studies in Science and Judaism, 7(??):285–318, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/221071.

Langermann:2008:MNG

[Lan08a] Y. Tzvi Langermann. From my notebooks: a (the?) geniza fragment on optics. Aleph: Historical Studies in Science and Judaism, 8(??):345–358, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/247249.

Langermann:2008:MTP

[Lan08b] Y. Tzvi Langermann. My truest perplexities. Aleph: Historical Studies in Science and Judaism, 8(??):301–317, ???? 2008. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/247253.

Langermann:2009:MNM

[Lan09] Y. Tzvi Langermann. From my notebooks: Medicine, mechanics and magic from Moses ben judah Galeano’s Ta‘alumot h.okmah. Aleph: Historical Studies in Science and Judaism, 9(2):353–377, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/317033.

Langermann:2010:MNS

[Lan10] Y. Tzvi Langermann. From my notebooks: Studies on the Hebrew Geminos: The chapter on weather signs. Aleph: Historical Studies in Science and Judaism, 10(2):357–395, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/394668.

Langermann:2011:MMM

[Lan11] Y. Tzvi Langermann. Materia medica et magica from animals, including a long, unknown passage from al–Mas‘ūdī. Aleph: Historical Studies in Science and Judaism, 11(1):169–178, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/423513.

Langermann:2012:NRN

[Lan12] Y. Tzvi Langermann. No reagent, no reaction: The barren transmission of Avicennan dynamics to H .asdai Crescas. Aleph: Historical Studies in Science and Judaism, 12(1):161–188, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/472830.

Langermann:2013:ATA

[Lan13a] Y. Tzvi Langermann. An alchemical treatise attributed to Joseph Solomon delmedigo. Aleph: Historical Studies in Science and Judaism, 13(1):77–94, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/505931.

Langton:2013:JRT

[Lan13b] Daniel R. Langton. Jewish religious thought, the Holocaust, and Darwinism: a comparison of Hans jonas and mordecai kaplan. Aleph: Historical Studies in Science and Judaism, 13(2):311–348, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/523713.

Langermann:2014:MNT

[Lan14] Y. Tzvi Langermann. From my notebooks: On Tajriba/Nissayon (“Experience”): Texts in Hebrew, Judeo-Arabic, and Arabic. Aleph: Historical Studies in Science and Judaism, 14(2):147–176, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/554176.

Langermann:2016:FCA

[Lan16] Y. Tzvi Langermann. Fragments of commentaries on Aristotle’s Physics from the David Kaufmann Genizah Collection, by Ibn Daud and others (?). Aleph: Historical Studies in Science and Judaism, 16(1):38–60, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/612501.

Langermann:2017:AB

[Lan17a] Y. Tzvi Langermann. Appendix B. Aleph: Historical Studies in Science and Judaism, 17(1):88–89, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/652311#sub02.

Langermann:2017:MNI

[Lan17b] Y. Tzvi Langermann. From my notebooks Isaac Israeli (the Elder): Some interesting remarks on the posterior analytics in his book on fevers. Aleph: Historical Studies in Science and Judaism, 17(1): 157–166, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
652314.

Langermann:2019:TJA

[Lan19] Y. Tzvi Langermann. Three judeo-Arabic marginalia on new materia medica from the new world and China. Aleph: Historical Studies in Science and Judaism, 19(1):137–155, ???? 2019. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/724814.

Lasker:2010:EEK

[Las10] Daniel J. Lasker. Eastern European Karaite attitudes towards modern science. Aleph: Historical Studies in Science and Judaism, 10(1):119–136, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/376442.

Lausch:2002:IHB

[Lau02] Hans Lausch. “The ignorant hold back their judgment and await the conclusions of the knowing”: Moses Mendelssohn and other mathematicians. Aleph: Historical Studies in Science and Judaism, 2(??):93–109, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407507/pdf.

Levy:2006:SMM

[LB06] Tony Levy and Charles Burnett. Sefer ha-Middot: a Mid-Twelfth-Century text on arithmetic and geometry attributed to Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 6(??): 57–238, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
203926.

Leicht:2013:RAM

[Lei13] Reimund Leicht. The reception of astrology in Medieval Ashkenazi culture. Aleph: Historical Studies in Science and Judaism, 13(2): 201–234, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
523712.

Levy:2001:HLV

[Lév01] Tony Lévy. Hebrew and Latin versions of an unknown mathematical text by Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 1(??):295–305, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/406599/pdf.

Levy:2002:NDP

[Lév02] Tony Lévy. A newly-discovered partial Hebrew version of al– Khwārizmī’s algebra. Aleph: Historical Studies in Science and Judaism, 2(??):223–234, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407513/pdf.

Levy:2012:IBJ

[Lév12] Tony Lévy. Immanuel ben Jacob of Tarascon (Fourteenth Century) and Archimedean geometry: an alternative proof for the area of a circle. Aleph: Historical Studies in Science and Judaism, 12(1):135–158, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/472829.

Leicht:2019:EF

[LF19] Reimund Leicht and Resianne Fontaine. Editors’ foreword. Aleph: Historical Studies in Science and Judaism, 19(1):7–8, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/724811.

Levy:2018:HAM

[LV18] Tony Lévy and Bernard Vitrac. Hero of Alexandria and mordekhai komtino: The encounter between mathematics in Hebrew and the Greek metrological corpus in Fifteenth-Century Constantinople. Aleph: Historical Studies in Science and Judaism, 18(2):181–262, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/706085.

Mancha:2002:LBG

[Man02] J. L. Mancha. Levi Ben Gerson’s star list for 1336. Aleph: Historical Studies in Science and Judaism, 2(??):31–57, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/407505/pdf.

Mancha:2012:TAR

[Man12] J. L. Mancha. The theory of access and recess in Levi ben Gerson’s astronomy, and its sources. Aleph: Historical Studies in Science and Judaism, 12(1):37–64, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/472826.

Mancha:2017:DLP

[Man17] José Luis Mancha. The dedicatory letter and the prologue of Levi ben Gershom’s: Tractatus instrumenti astronomiae. Aleph: Historical Studies in Science and Judaism, 17(2):220–249, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/669422.

Moreno-Carvalho:2002:NDL

[MC02] Francisco Moreno-Carvalho. A newly discovered letter by Galileo Galilei: Contacts between Galileo and Jacob Rosales (Manoel Bocarro Francês), a Seventeenth-Century Jewish scientist and Sebastianist. Aleph: Historical Studies in Science and Judaism, 2(??): 59–91, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
407506/pdf.

Mentgen:2009:JAC

[Men09] Gerd Mentgen. Jewish alchemists in central Europe in the later middle ages: Some new sources. Aleph: Historical Studies in Science and Judaism, 9(2):345–352, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/317026.

Mesler:2014:MLT

[Mes14] Katelyn Mesler. The medieval lapidary of Techel/Azareus on engraved stones and its Jewish appropriations. Aleph: Historical Studies in Science and Judaism, 14(2):75–143, ???? 2014. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/554175/pdf.

Meyrav:2022:AEA

[Mey22] Yoav Meyrav. Averroes’s epitome of Aristotle’s Physics in Hebrew: Translation, transmission, and revision. Aleph: Historical Studies in Science and Judaism, 22(1–2):183–233, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/874965.

Mancha:2005:LBG

[MF05] Jose Luis Mancha and Gad Freudenthal. Levi ben Gershom’s criticism of Ptolemy’s Astronomy. Aleph: Historical Studies in Science and Judaism, 5(??):35–167, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/187942.

Morrison:2011:ATM

[Mor11] Robert Morrison. An astronomical treatise by Mūsā Jālīnūs alias Moses galeano. Aleph: Historical Studies in Science and Judaism, 11(2):385–413, ???? 2011. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/453378.

North:2002:SJC

[Nor02] J. D. North. Some Jewish contributions to Iberian astronomy. Aleph: Historical Studies in Science and Judaism, 2(??):271–278, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407516/
pdf.

Pear:2015:DDA

[Pea15] Rachel S. A. Pear. Differences over Darwinism: American orthodox Jewish responses to evolution in the 1920s. Aleph: Historical Studies in Science and Judaism, 15(2):343–387, ???? 2015. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/590706.

Rodriguez-Arribas:2020:JBS

[RAK20] Josefina Rodríguez-Arribas and Maud Kozodoy. Joseph ben Solomon T .ait.as.aq on the construction of an astrolabe: Study, diplomatic edition, and annotated translation. Aleph: Historical Studies in Science and Judaism, 20(1–2):7–101, ???? 2020. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/776497.

Rand:2009:CRU

[Ran09] Michael Rand. Clouds, rain, and the upper Waters: From bereshit rabbah to the Piyyut.im of eleazar be-rabbi qillir. Aleph: Historical Studies in Science and Judaism, 9(1):13–39, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/260442.

Ravitzky:2008:AMM

[Rav08] Aviezer Ravitzky. Aristotle’s meteorology and the maimonidean modes of interpreting the account of creation. Aleph: Historical Studies in Science and Judaism, 8(??):361–400, ???? 2008. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/247256.

Reisman:2003:MFR

[Rei03] David C. Reisman. In memoriam: Franz Rosenthal August 31, 1914–April 8, 2003. Aleph: Historical Studies in Science and Judaism, 3(??):329–342, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/407530/pdf.

Reisman:2007:JRN

[Rei07] Arnold Reisman. Jewish refugees from Nazism, Albert Einstein, and the modernization of higher education in Turkey (1933–1945). Aleph: Historical Studies in Science and Judaism, 7(??):253–281, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/221073.

Rosenberg:2015:SRT

[RF15] Shalom Rosenberg and Gad Freudenthal. Science and religion (Torah u-Madda) in modern Jewish thought: an overview. Aleph: Historical Studies in Science and Judaism, 15(1):65–111, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/576358.

Robinson:2003:FRH

[Rob03] James T. Robinson. The first references in Hebrew to al–Bit.rūjī’s on the principles of astronomy. Aleph: Historical Studies in Science and Judaism, 3(??):145–163, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/407524/pdf.

Robinson:2009:SIT

[Rob09] James T. Robinson. Samuel Ibn Tibbon’s Peruš ha-Millot ha-Zarot and al–Fārābī’s Eisagoge and Categories. Aleph: Historical Studies in Science and Judaism, 9(1):41–76, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https:

//muse.jhu.edu/pub/3/article/260443.

Robinson:2021:SHS

[Rob21] James Theodore Robinson. The “Secret of the Heavens” and the “Secret of Number”: Immanuel of Rome’s mathematical supercommentaries on Abraham Ibn Ezra in his commentary on Qohelet 5:7 and 7:27. Aleph: Historical Studies in Science and Judaism, 21(2):279–308, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/845999.

Rubin:2008:UUA

[Rub08] Rehav Rubin. H .ug ha-’ares. by Rabbi Solomon of Chelm: an early geographical treatise and its sources. Aleph: Historical Studies in Science and Judaism, 8(??):131–147, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247257.

Ruderman:2002:SJR

[Rud02] David B. Ruderman. Some Jewish responses to smallpox prevention in the Late Eighteenth and Early Nineteenth Centuries: a new perspective on the modernization of European Jewry. Aleph: Historical Studies in Science and Judaism, 2(??):111–144, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/407508/pdf.

Schwartz:2001:TME

[Sch01] Dov Schwartz. From theurgy to magic: The evolution of the magical-talismanic justification of sacrifice in the circle of nahmanides and his interpreters. Aleph: Historical Studies in Science and Judaism, 1(??):165–213, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/406596/pdf.

Schwartz:2003:CAM

[Sch03] Dov Schwartz. Conceptions of astral magic within Jewish rationalism in the Byzantine empire. Aleph: Historical Studies in Science and Judaism, 3(??):165–211, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/407525/pdf.

Sela:2001:FBB

[Sel01] Shlomo Sela. The fuzzy borders between astronomy and astrology in the thought and work of three Twelfth-Century Jewish intellectuals. Aleph: Historical Studies in Science and Judaism, 1(??): 59–100, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
406593/pdf.

Sela:2004:QAS

[Sel04] Shlomo Sela. Queries on astrology sent from southern France to Maimonides: Critical edition of the Hebrew text, translation, and commentary. Aleph: Historical Studies in Science and Judaism, 4(??):89–190, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/176046.

Sela:2009:STU

[Sel09] Shlomo Sela. Sefer ha-Tequfah: an unknown treatise on anniversary horoscopy by Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 9(2):240–254, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/317025.

Sela:2010:FUR

[Sel10] Shlomo Sela. A fragment from an unknown redaction of Re’šit h.ok̄mah by Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 10(1):43–66, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/376440.

Sela:2012:E

[Sel12a] Shlomo Sela. Erratum. Aleph: Historical Studies in Science and Judaism, 12(2):413, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/487985.

Sela:2012:MMN

[Sel12b] Shlomo Sela. Maimonides and Māshā’allāh on the ninth orb of the signs and astrology. Aleph: Historical Studies in Science and Judaism, 12(1):101–134, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/472828.

Sela:2013:NIE

[Sel13] Shlomo Sela. A newly identified essay on anniversary horoscopy embedded in Abraham Bar H .iyya’s astronomical tables: Hebrew edition, translation and commentary. Aleph: Historical Studies in Science and Judaism, 13(1):27–76, ???? 2013. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/505930.

Sela:2014:AAE

[Sel14] Shlomo Sela. The astrological-astronomical encyclopedia in MS Paris 1058. Aleph: Historical Studies in Science and Judaism, 14(1):189–241, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.edu/

pub/3/article/543889.

Sela:2016:AFP

[Sel16] Shlomo Sela. Al-farghānī on the 48 Ptolemaic constellations: a newly discovered text in Hebrew translation. Aleph: Historical Studies in Science and Judaism, 16(2):249–365, ???? 2016. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/627990.

Sela:2017:GAA

[Sel17] Shlomo Sela. Gersonides’ astrology and Abraham Ibn Ezra. Aleph: Historical Studies in Science and Judaism, 17(2):250–323, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/669423.

Sela:2019:AIE

[Sel19] Shlomo Sela. The Abraham Ibn Ezra–Peter of Limoges astrological–exegetical connection. Aleph: Historical Studies in Science and Judaism, 19(1):9–57, ???? 2019. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/724815.

Sela:2020:HBT

[Sel20] Shlomo Sela. Henry Bate, translator of Abraham Ibn Ezra’s astrological writings. Aleph: Historical Studies in Science and Judaism, 20(1–2):103–201, ???? 2020. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/776498.

Sela:2022:AIE

[Sel22a] Shlomo Sela. Abraham Ibn Ezra’s scholarly writings: an update. Aleph: Historical Studies in Science and Judaism, 22(1–2):287– 300, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
874968.

Sela:2022:HUT

[Sel22b] Shlomo Sela. A hitherto unknown translation by Abraham Ibn Ezra of a treatise by Abū Ma‘shar: Edition of the French translation by Hagin le Juif and English translation. Aleph: Historical Studies in Science and Judaism, 22(1–2):9–38, ???? 2022. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/874960.

Sela:2006:AIE

[SF06] Shlomo Sela and Gad Freudenthal. Abraham Ibn Ezra’s scholarly writings: a chronological listing. Aleph: Historical Studies in Science and Judaism, 6(??):13–55, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.

jhu.edu/pub/3/article/203927.

Shapira:2018:LJF

[Sha18] Dan Shapira. The Logika of the Judaizers. A Fifteenth-Century Ruthenian Translation from Hebrew. Critical Edition of the Slavic Texts Presented alongside their Hebrew Sources with Introduction, English Translation, and Commentary ed. by Dan Shapira (review). Aleph: Historical Studies in Science and Judaism, 18(2): 293–310, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
706091.

Shiloah:2004:MCW

[Shi04] Amnon Shiloah. Musical concepts in the works of saadia gaon. Aleph: Historical Studies in Science and Judaism, 4(??):265–282, ???? 2004. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/176047.

Shulman:2002:TIC

[Shu02] David Shulman. Is there an Indian connection to Sefer yes.irah? Aleph: Historical Studies in Science and Judaism, 2(??):191–199, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/407511/
pdf.

Stern:2015:ACS

[SI15] Sacha Stern and Justine Isserles. The astrological and calendar section of the earliest Mah.zor Vitry manuscript (MS ex-Sassoon 535). Aleph: Historical Studies in Science and Judaism, 15(2): 199–317, ???? 2015. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
590707.

Szilagyi:2016:FCP

[SL16] Krisztina Szilágyi and Y. Tzvi Langermann. A fragment of a composition on physics by Abraham Ibn Daud in Judeo–Arabic: an edition of the text. Aleph: Historical Studies in Science and Judaism, 16(1):32–38, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/612500.

Sladek:2022:DGM

[Slá22] Pavel Sládek. David Gans’s Magen David: Text and context. Aleph: Historical Studies in Science and Judaism, 22(1–2):235– 276, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
874966.

Stern:2007:ATS

[SM07] Sacha Stern and Piergabriele Mancuso. An astronomical table by Shabbetai Donnolo and the Jewish calendar in Tenth-Century Italy. Aleph: Historical Studies in Science and Judaism, 7(??): 13–41, ???? 2007. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
221074.

Stern:2008:SDJ

[SM08] Sacha Stern and Piergabriele Mancuso. Shabbetai donnolo and the Jewish calendar: Corrigenda. Aleph: Historical Studies in Science and Judaism, 8(??):343–344, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247250.

Smithuis:2006:AIE

[Smi06] Renate Smithuis. Abraham Ibn Ezra’s astrological works in Hebrew and Latin: New discoveries and exhaustive listing. Aleph: Historical Studies in Science and Judaism, 6(??):239–338, ???? 2006. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/203928.

Sela:2009:THF

[SS09] Shlomo Sela and Renate Smithuis. Two Hebrew fragments from unknown redactions of Abraham Ibn Ezra’s Sefer ha–Mivh.arim and Sefer ha-Še’elot. Aleph: Historical Studies in Science and Judaism, 9(2):225–240, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/317027.

Stern:2005:REA

[Ste05] Sacha Stern. Rachel elior on ancient Jewish calendars: a critique. Aleph: Historical Studies in Science and Judaism, 5(??):287–292, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/187944.

Stern:2008:KNW

[Ste08] Josef Stern. The knot that never was. Aleph: Historical Studies in Science and Judaism, 8(??):319–339, ???? 2008. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/247252.

Stroumsa:2001:RMC

[Str01] Sarah Stroumsa. “Ravings”: Maimonides’ concept of pseudoscience. Aleph: Historical Studies in Science and Judaism, 1(??): 141–163, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
406595/pdf.

Stroumsa:2014:ML

[Str14] Sarah Stroumsa. On Maimonides and on logic. Aleph: Historical Studies in Science and Judaism, 14(1):259–263, ???? 2014. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/543891.

Szilagyi:2016:FBP

[Szi16] Krisztina Szilágyi. A fragment of a book of physics from the David Kaufmann Genizah Collection (Budapest) and the identity of Ibn Daud with Avendauth. Aleph: Historical Studies in Science and Judaism, 16(1):10–31, ???? 2016. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/612499.

Taube:2010:TST

[Tau10] Moshe Taube. Transmission of scientific texts in 15th-century Eastern knaan. Aleph: Historical Studies in Science and Judaism, 10(2):315–353, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/394667.

Vallicrosa:2010:PAC

[Val10] José María Millás Vallicrosa. Pedro Alfonso’s contribution to astronomy. Aleph: Historical Studies in Science and Judaism, 10(1): 139–168, ???? 2010. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
376443.

Vidro:2017:OYCa

[Vid17a] Nadia Vidro. The origins of the 247-year calendar cycle. Aleph: Historical Studies in Science and Judaism, 17(1):95–137, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/652312.

See erratum [Vid17b].

Vidro:2017:OYCb

[Vid17b] Nadia Vidro. The origins of the 247-year calendar cycle — erratum. Aleph: Historical Studies in Science and Judaism, 17(2):371, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/669426.

See [Vid17a].

Vidro:2021:NRJ

[Vid21] Nadia Vidro. Non-rabbanite Jewish calendars in the works of Jacob al–Qirqisānī and Saadia Gaon. Aleph: Historical Studies in Science and Judaism, 21(1):149–187, ???? 2021. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/799681.

Vidro:2022:MQC

[Vid22] Nadia Vidro. The Medieval Qaraite calendar in the Diaspora. Aleph: Historical Studies in Science and Judaism, 22(1–2):125– 156, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
874963.

Visi:2014:BBN

[Vis14] Tamás Visi. Berechiah ben Nat.ronai ha-Naqdan’s Dodi ve–Nek̄di and the transfer of scientific knowledge from Latin to Hebrew in the Twelfth Century. Aleph: Historical Studies in Science and Judaism, 14(2):9–75, ???? 2014. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/554174/pdf.

Visi:2021:MIC

[Vis21] Tamás Visi. The “Meteorological” interpretation of the creation narrative from John Philoponus to Saadia Gaon. Aleph: Historical Studies in Science and Judaism, 21(2):209–278, ???? 2021. CO-DEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL

https://muse.jhu.edu/pub/3/article/846002.

Visi:2022:MIC

[Vis22] Tamás Visi. The “Meteorological” interpretation of the creation narrative: John Philoponus’s legacy in Abraham Ibn Ezra and Moses Maimonides. Aleph: Historical Studies in Science and Judaism, 22(1–2):39–99, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/874961.

Volkov:2001:JSI

[Vol01] Shulamit Volkov. Jewish scientists in Imperial Germany (parts I and II). Aleph: Historical Studies in Science and Judaism, 1(??): 215–216, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
406597/pdf.

Wasserstrom:2002:FTO

[Was02] Steven M. Wasserstrom. Further thoughts on the origins of Sefer yes.irah. Aleph: Historical Studies in Science and Judaism, 2(??): 201–221, ???? 2002. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
407512/pdf.

Xiuyuan:2022:MCP

[Xiu22] Dong Xiuyuan. Maimonides’ cosmogony-prophetology puzzle: Revisiting the traditionalist approach. Aleph: Historical Studies in Science and Judaism, 22(1–2):101–123, ???? 2022. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/874962.

Yoeli-Tlalim:2018:EPL

[YT18] Ronit Yoeli-Tlalim. Exploring Persian lore in the Hebrew book of asaf. Aleph: Historical Studies in Science and Judaism, 18(1): 123–145, ???? 2018. CODEN ???? ISSN 1553-3956 (print), 1565-

1525 (electronic). URL https://muse.jhu.edu/pub/3/article/
693216.

Zalkin:2005:SLC

[Zal05] Mordekhai Zalkin. Scientific literature and cultural transformation in Nineteenth-Century East European Jewish society. Aleph: Historical Studies in Science and Judaism, 5(??):249–271, ???? 2005. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (elec-

tronic). URL https://muse.jhu.edu/pub/3/article/187945.

Zonta:2009:NGS

[ZF09] Mauro Zonta and Gad Freudenthal. Nicomachus of Gerasa in Spain, circa 1100: Abraham Bar H .iyya’s testimony. Aleph: Historical Studies in Science and Judaism, 9(2):189–224, ???? 2009. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/317024.

Zonta:2001:NDJ

[Zon01] Mauro Zonta. New data on Judah Messer Leon’s commentaries on the physics. Aleph: Historical Studies in Science and Judaism, 1(??):307–323, ???? 2001. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/406600/pdf.

Zonta:2003:HTH

[Zon03] Mauro Zonta. A Hebrew translation of Hippocrates’ De superfoetatione: Historical introduction and critical edition. Aleph: Historical Studies in Science and Judaism, 3(??):97–113, ???? 2003. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic).

URL https://muse.jhu.edu/pub/3/article/407523/pdf.

Zonta:2012:TIS

[Zon12] Mauro Zonta. Two implicit sources of Judah Messer Leon’s commentary on physics 1–4. Aleph: Historical Studies in Science and Judaism, 12(1):189–208, ???? 2012. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://muse.jhu.

edu/pub/3/article/472831.

Zonta:2017:TNH

[Zon17] Mauro Zonta. Twenty-nine Hebrew glosses on Averroes’ middle commentary on Aristotle’s Metaphysics. Aleph: Historical Studies in Science and Judaism, 17(2):335–353, ???? 2017. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/669424.

Zuber:2020:MIF

[ZP20] Mike A. Zuber and Rafał T. Prinke. Mistaken identity, forged prophecies, and a Paracelsus commentary: The elusive Jewish alchemist mardochaeus de nelle. Aleph: Historical Studies in Science and Judaism, 20(1–2):263–299, ???? 2020. CODEN ???? ISSN 1553-3956 (print), 1565-1525 (electronic). URL https://

muse.jhu.edu/pub/3/article/776499.